OpenStudy (anonymous):

Factor completely: 2x2 + 4x − 2 2(x2 + 2x − 1) 2(x2+ 2) 2x(x2 + 2x − 1) Prime

OpenStudy (anonymous):

a

OpenStudy (anonymous):

\[2 x ^{2} + 4x - 2\] \[2(x ^{2} + 2x - 1)\] this cannot be factored any further

OpenStudy (anonymous):

\[2x ^{2}+4x-2=2\left( x ^{2}+2x-1 \right)\] \[=2\left( x ^{2}+2x+1-1-1 \right) \] \[=2\left[ \left( x+1 \right)^{2}-\left( \sqrt{2} \right)^{2^{}} \right]\] \[=2\left( x+1+\sqrt{2} \right)\left( x+1-\sqrt{2} \right)\]